OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. A(4, 1), B(5, 2)

OpenStudy (cp9454):

y=x-2

OpenStudy (anonymous):

how?

OpenStudy (cp9454):

find slope \[\frac{ 2-1 }{ 5-4 }\]

OpenStudy (anonymous):

i understand that, but i don't know where to go from there

OpenStudy (cp9454):

then apply slope point formula \[(y-y _{1}) = m(x-x _{1})\]

OpenStudy (cp9454):

where \[(x _{1} , x _{2}) = (4 , 1)\]

OpenStudy (anonymous):

i'm still not sure... where did you get the 4, 1? just from the original 2 points or what

OpenStudy (cp9454):

ya that,s right . you can go through a book it is online download it and you will understand everything.

OpenStudy (anonymous):

listen to him books help with everthing